Can you help me formulate an equation for the following homework problem? It did not come from a book. Thanks
Two friends live 15 miles apart. One day, they decide to jog and meet each other. Tanya leaves her house and heads east, jogging at a rate of 2.5 mph. At the same time Kelly leaves her house and heads west, jogging at 3.5 mph. How long does it take for the two to meet?
15= 2.5x for Tanya and
15= 3.5 x for Kelly? This can't be right, as I can only have one x. Please help

Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

Distance=rate*time or time=distance/rate & seeing as the rates are accumulative, we have:
15/(2.5+3.5)=x now cross multiply.
6x=15
x=2.5 hours for them to meet.
Proof
2.5*2.5+3.5*2.5=15
6.25+8.75=15
15=15


Answer by stanbon(57347)

Two friends live 15 miles apart. One day, they decide to jog and meet each other. Tanya leaves her house and heads east, jogging at a rate of 2.5 mph. At the same time Kelly leaves her house and heads west, jogging at 3.5 mph. How long does it take for the two to meet?
------------------------
They are jogging toward one another so neither needs to jog 15 miles.
---------------------
Tanya DATA:
Rate = 2.5 mph; distance = x miles; time = d/r = x/2.5 hrs
----------------
Kelly DATA:
Rate = 3.5 mph; distance = "15-x" miles ; time = (15-x)/3.5 hrs
-------------------------
Tanya time = Kelly time
x/2.5 = (15-x)/3.5
3.5x = 37.5-2.5x
6x = 37.5
x = 6.25 miles
----------------
Tanya time = x/2.5 = 6.25/2.5 = 2.5 hrs
--------------
Kelly time = (15-x)/3.5 = 8.75/3.5 = 2.5 hrs
===================
Cheers,
Stan H.